For the following circuit, find the following: Wh
Answer #1

a)
1/R1 = 1/3 + 1/6
R1 = 18/9 = 2 ohms

1/R2 = 1/4 + 1/12
R2 = 48/16
R2 = 3 Ohms

Net resistance = R1 + R2 = 5 Ohms

20 = 5 * I
I = 4 A

b) Potential Difference across 3 and 6 ohms :
Hence
V1 = 4*2 = 8 V

Potential Difference across 4 and 12 ohms :
V2 = 4*3 = 12 V

c) P = I^2 x R
Plug in the values and calculate